# Better Late Than Never: Formulating and Benchmarking Recommendation Editing

### v1

@@ 3 The Setup of Recommendation Editing
4 Experimental Results
Table 2: Editing effectiveness of all editing methods. The table presents the reported values for the indicators ES, EC, EP, and EA, expressed as percentages. The best results are in bold, and the second-best results are underlined.
@@ 4 Experimental Results
4.1 Editing Effectiveness
@@
  • Graph is not necessary in recommendation editing. The EFT method shows superior editing performance in graph-based collaborative filtering methods (LightGCN, XSimGCL), differing from the basic Fine-Tune method only in graph visibility. While graph knowledge usually benefits collaborative filtering, in recommendation editing, the reliance on positive feedback introduces discrepancies with the negative feedback targeted for edits, potentially degrading performance.
  • The regularization-based approach improves recommendation editing performance by enhancing Editing Prudence. The regularization-based approach shows suboptimal performance in various recommendation editing tasks. Unlike basic fine-tuning, which lacks constraints, this method uses regularization to minimize divergence from the original model, potentially improving editing caution. However, it can hinder collaborative editing and overly strict constraints may reduce editing accuracy. For example, the L2 method only achieved an editing accuracy of 0.91 on the XSimGCL model using the QB-video dataset, while ideal accuracy approaches 1.
  • The replay-based approach improves recommendation editing performance by enhancing Editing Collaboration. Contrary to regularization, the replay-based approach boosts recommendation editing by enhancing collaboration, albeit at the expense of prudence. SPMF and RSR excel in editing MF models but are less effective in graph-based models compared to basic fine-tuning. This approach compensates for missing collaborative information in MF models but magnifies inaccuracies in graph-based models. Notably, SPMF outperforms RSR, suggesting a potential for developing more effective replay strategies for recommendation editing.
  • Existing recommendation system methods are not suitable for recommendation editing. We evaluated several recommendation system methods such as SML, SRIU, SiReN, and SiGRec for their compatibility with recommendation editing objectives. Unfortunately, they underperformed, significantly lagging behind basic fine-tuning. This gap likely stems from a mismatch between the methods’ original design goals and the specific demands of recommendation editing.
@@
    4.2 Editing Efficiency
@@
Figure 2: The impact of the number of explicit editing pairs on editing effectiveness.
Figure 3: Editing efficiency.
Figure 4: The influence of proper objective.
4.3 Explicit Editing Pairs Quantity
In real-world recommendation editing, the number of explicit editing pairs is often uncertain and impacts recommendation editing differently. We assess this effect on recommendation editing performance metrics. Figure 2 shows how the ES, EC, EP, and EA metrics fluctuate with varying numbers of explicit editing pairs, particularly when editing the XSimGCL model on the QB-video dataset. The EFT method consistently excels in ES, while regularization-based and replay-based approaches maintain stability in ES. Conversely, methods like SML and EGNN exhibit worsening performance with more editing pairs. For EC and EP, most methods show an increase in EC and a decrease in EP with more pairs, due to the additional collaborative information complicating the editing task, thus degrading EP performance. For EA, EFT maintains 100% accuracy, whereas other methods decline as editing pairs increase. The L2 method uniquely shows opposite trends in EC and EP and suffers the most in EA, likely due to its overly stringent constraints.
4.4 Editing Objective
@@
5 Related Works
Online/Incremental Recommendation. With the exponential growth of online information, RS typically operates incrementally, continuously adapting to new users, items, and interactions Chen et al. (2024); He et al. (2020); Koren et al. (2021); Wang et al. (2019). Recent efforts have enhanced efficient model updates for online learning in RS. Continual learning approaches, such as ADER Mi et al. (2020), utilize adaptive distillation loss to periodically integrate feedback by replaying past training samples. ReLoop Cai et al. (2022) introduces a self-correction loop with a custom loss based on user feedback. Incremental learning strategies aim for rapid adaptation to new data with minimal overhead. MAN Mi and Faltings (2021) updates a neural RS using a continuously updated nonparametric memory, while FIRE Xia et al. (2022) employs a fast incremental learning method from a graph signal processing viewpoint. Online/incremental recommendation systems continually update models based on new user feedback, yet they significantly differ from recommendation editing. Although they process negative feedback and aim for broad accuracy, they do not specifically correct erroneous recommendations. This often results in compromises in addressing these errors due to the generalized modeling of extensive user data. In contrast, recommendation editing focuses on precisely correcting specific errors, fundamentally differing from the broader adaptation objectives of online methods.
Recommendation with Negative Feedback. Most recommendation systems primarily focus on users’ positive historical feedback, often neglecting negative feedback such as low ratings, content skips, and dislikes Huang et al. (2023). Integrating negative feedback has proven to significantly enhance system performance Jeunen (2019). To leverage this valuable information, some approaches now blend both positive and negative feedback. For instance, movie recommendation systems Chen et al. (2021) construct separate user profiles for positive and negative feedback to improve suggestions. Similarly, SiReN Seo et al. (2022) and SiGRec Huang et al. (2023) employ Signed Graphs to incorporate both feedback types, boosting performance with graph-based methods. However, these methods, which balance positive and negative inputs to model user preferences, require training from scratch and are often impractical for recommendation editing due to the need for complete historical feedback data and the prevalence of incomplete datasets.
Model Editing. Model editing, initially introduced within the context of generative adversarial networks Bau et al. (2020), involves modifying neural network layers—treated as linear associative memories—to update weights based on specified inputs or concepts. Emerging alternative methods include training hyper-networks for weight adjustments De Cao et al. (2021); Mitchell et al. (2021), surgically altering “knowledge neurons”Dai et al. (2022), and using auxiliary detectors to redirect relevant inputs to modified modelsMitchell et al. (2022). Model editing has also been applied to open vocabulary image-language models like CLIP Radford et al. (2021), demonstrating how weight interpolation can balance performance between original and edited tasks Ilharco et al. (2022). Despite its prominence in computer vision and natural language processing, model editing does not fully address the distinct challenges in recommendation systems, which deal with non-independent and identically distributed (non-iid) data and dynamic user-item interactions. Unlike conventional tasks like classification that focus on individual instances, recommendation editing emphasizes ranking, necessitating specialized approaches to effectively manage the unique ranking challenges specific to recommendation systems.
6 Conclusion
@@ Appendix A Experimental Setting
A.1 Recommendation Editing Setting
In our experiments, we randomly split the positive feedback in the dataset into a training set and a test set at a ratio of 8:2. We use the training set to train the initial collaborative filtering model. For the same dataset and backbone, all editing methods underwent testing on the same initial model, The performance before editing is documented in Appendix B.1. We utilized the full set of negative feedback from the dataset as expected editing pairs, $\mathcal{E}$, randomly selected 10 user-item pairs from the intersection of the initial model’s top 50 recommendation results and the negative feedback to serve as explicit editing pairs, $\mathcal{E}^{\mathbb{E}}$. The remaining negative feedback pairs were considered implicit editing pairs, $\mathcal{E}^{\mathbb{I}}$, since the similarity pairs of $\mathcal{E}^{\mathbb{E}}$ are challenging to define strictly. Pairs recommended by the initial model that did not involve negative feedback were defined as unnecessary editing pairs, $\bar{\mathcal{E}}$. Apart from RSR, SPMF, SML, and SiGRec, which perform editing through their unique designs, all other methods employ E-BPR Loss as optimization objectives for editing. To ensure the efficiency of model editing, all finetuning-based methods are restricted to a maximum of 20 rounds of fine-tuning. We evaluated the editing performance of all methods through ES (editing score), EC (editing collaboration), EP (editing prudence), EA (editing accuracy), and Recall and NDCG metrics in the test set after model editing. All methods were repeated 10 times, and the average performance was reported.
A.2 Dataset
Epinions Zhao et al. [2014], sourced from online consumer review platform Epinions.com, comprises user ratings with a range from 1 to 5, detailed item reviews, and directed trust relationships among users. We classify interactions as positive feedback if ratings exceed 3, and negative feedback if ratings are 3 or lower.
KuaiRand Gao et al. [2022], an unbiased sequential recommendation dataset, is derived from the recommendation logs of Kuaishou, a popular video-sharing mobile application. A user clicking on a video indicates positive feedback, while negative feedback is represented by items shown to a user who then clicks the dislike button.
QB-video Yuan et al. [2022] is originated from the QQ browser. The positive feedback here is represented by the user clicking on videos, and negative feedback occurs when a user is shown a video but does not click on it.
A.3 Collaborative filtering model
@@
MF Rendle et al. [2009a] is a cornerstone in collaborative filtering for recommendation systems, decomposing a user-item interaction matrix into lower-dimensional latent factor matrices for users and items. This approach uncovers hidden patterns in data, enabling predictions of unobserved interactions by capturing underlying user preferences and item characteristics.
LightGCN He et al. [2020] simplifies the design of Graph Convolutional Networks (GCNs) for collaborative filtering in recommendation systems. It streamlines GCNs by removing feature transformation and nonlinear activation, focusing on the essential component of neighborhood aggregation to learn user and item embeddings. This process effectively captures the high-order connectivity in the user-item interaction graph, enhancing recommendation quality by leveraging the structural information within the graph.
XSimGCL Yu et al. [2023] is an improvement upon SimGCL, introducing a highly straightforward graph contrastive learning approach. Not only does it inherit the noise-based augmentation from SimGCL, but it also significantly reduces computational complexity by simplifying the propagation process.
A.4 Model Parameter Settings
@@
A.5 Editing Method

### v2

@@ 5.1.1. Datasets
Table 2. Editing effectiveness of all editing methods. The table presents the reported values for the indicators ES, EC, and EP, expressed as percentages. The best results are in bold, and the second-best results are underlined.Different types of methods are highlighted with different colors(Regularization-based, Replay-based, Optimization-based, CF with Negative Feedback, Model Editing and Finetuning-based).
@@
5.1.2. Editing Methods
To evaluate the performance of different methods, we categorize various methods into six categories associated with incremental learning and model editing:
Finetuning-based Approaches: FT (Fine-Tuning) (Sinitsin et al., 2020) and EFT (Editing Fine-Tuning). These methods edit the model using a fine-tuning manner based on explicit editing data. For example, the FT method modifies the recommendation model by changing its original parameters, requiring the entire graph’s information for graph-based collaborative filtering models. In contrast, EFT modifies the final user and item embeddings through a model that generates these embeddings, without requiring detailed knowledge of the model’s architecture and training data.
Regularization-based Approaches: LWF (Learning Without Forgetting) (Li and Hoiem, 2017), L2 (Online L2Reg) (Lin et al., 2022), SRIU (Sample Reweighting Incremental Update) (Peng et al., 2021). These methods also edit models by fine-tuning explicit editing data and using the original model for regularization during the editing process. LWF normalizes the output on the original model using editing data, and L2 utilizes the parameters of the original model for regularization. SRIU contends that overfitting editing data with low confidence in the original model can lead the edited model to deviate from the original model’s patterns. It adopts a re-weighting approach that assigns lower weights to data with low confidence.
Replay-based Approaches: RSR (Random Sampling with a Reservoir) (Vitter, 1985) and SPMF (Stream-centered Probabilistic Matrix Factorization model) (Wang et al., 2018). This category method balances the efficiency and effectiveness of model editing by sampling a small number of training examples. The model is edited by fine-tuning based on both the sampled data and the explicit editing data. RSR adopts random sampling, while SPMF uses the original model’s prediction scores on training samples to perform weighted sampling.
Optimization-based Approaches: SML (Sequential
Meta-Learning) (Zhang et al., 2020). SML edits the model by training a Transfer module to generate the new parameters after model editing. Due to the constraints of our task, we cannot train SML on sequential data as originally designed for SML. Instead, we simulate editing data by randomly sampling the recommendation results of the original model multiple times. This simulated data is then used to train the Transfer module for the editing process.
CF with Negative Feedback: SiReN (Sign-aware Recommender system based on GNN models) (Seo et al., 2022) and SiGRec (Signed Graph Neural Network Recommendation model) (Huang et al., 2023). These methods handle both positive and negative feedback during collaborative filtering model training. In this paper, we adopt these approaches to handle negative feedback for model editing purposes.
Editing Approach: EGNN (Editable Graph Neural Networks) (Liu et al., 2023) and BiEGNN proposed by ourselves. EGNN enhances editing capabilities through an additional MLP (Multi-Layer Perceptron) module. Given the lack of node attributes in collaborative filtering tasks, we use the representations generated by the original model as inputs for the editing module. The existence of two types of nodes (i.e., user and item) utilizing the same editing module may create bias. To tackle this issue, we propose an improved version of EGNN, named BiEGNN, which trains distinct editing modules for users and items respectively to conduct editing.
To verify the generalization of editing methods, we conducted experiments on three representative recommendation models MF(Matrix Factorization) (Rendle et al., 2009), LightGCN (He et al., 2020) and XSimGCL (eXtremely Simple Graph Contrastive Learning) (Yu et al., 2023) simultaneously.
5.1.3. Model Editing Setting
In our experiments, we randomly split the positive feedback in the dataset into a training set and a test set at a ratio of 8:2 (He et al., 2020). We use the training set to train the initial collaborative filtering model. For the same dataset and backbone, all editing methods underwent testing on the same initial model. We used the full set of negative feedback from the dataset as expected editing pairs $\mathcal{E}$, randomly sampling 10 user-item pairs from the intersection of the initial model’s recommendation results. The editing pairs consist of the top 50 items predicted for each user and the correct editing pairs to form explicit editing pairs $\mathcal{E}^{\mathbb{E}}$. Apart from RSR, SPMF, SML, and SiGRec, which perform editing through their unique designs, all other methods employ E-BPR Loss as optimization objectives for editing. To ensure the efficiency of model editing, all finetuning-based methods are restricted to a maximum of 20 rounds of fine-tuning. We evaluated the editing performance of all methods through ES (editing score), EC (editing collaboration), EP (editing prudence), EA (editing accuracy), and Recall and NDCG metrics in the test set after model editing. All methods were repeated 10 times, and the average performance was reported.
Figure 2. Editing efficiency.
Figure 3. The impact of the number of explicit editing pairs on editing effectiveness.
5.2. Editing Effectiveness (RQ1)
@@
  • Graph is not necessary in model editing. The EFT method shows superior editing performance in graph-based collaborative filtering methods (LightGCN, XSimGCL), differing from the basic Fine-Tune method only in graph visibility. While graph knowledge usually benefits collaborative filtering, in recommendation editing, the reliance on positive feedback introduces discrepancies with the negative feedback targeted for edits, potentially degrading performance.
  • The regularization-based approach improves model editing performance by enhancing Editing Prudence. The regularization based approach shows suboptimal performance in various recommendation editing tasks. Unlike basic fine-tuning, which lacks constraints, this method uses regularization to minimize divergence from the original model, potentially improving editing caution. However, it can hinder collaborative editing and overly strict constraints may reduce editing accuracy. For example, the L2 method only achieved an editing accuracy of 0.91 on the XSimGCL model using the QB-video dataset, while ideal accuracy approaches 1.00.
  • The replay-based approach improves model editing performance by enhancing Editing Collaboration. Contrary to regularization, the replay-based approach boosts recommendation editing by enhancing collaboration, albeit at the expense of prudence. SPMF and RSR excel in editing MF models but are less effective in graph-based models compared to basic fine-tuning. This approach compensates for missing collaborative information in MF models but magnifies inaccuracies in graph-based models. Notably, SPMF outperforms RSR, suggesting a potential for developing more effective replay strategies for recommendation editing.
  • Existing recommendation system methods are not suitable for model editing. We evaluated several recommendation system methods such as SML, SRIU, SiReN, and SiGRec for their compatibility with recommendation editing objectives. Unfortunately, they underperformed, significantly lagging behind basic fine-tuning. This gap likely stems from a mismatch between the methods’ original design goals and the specific demands of recommendation editing.
@@
  • The effectiveness of editing is influenced by the intensity of negative feedback. From the perspective of different datasets, we can observe that most Editing Methods perform weaker on the KuaiRand dataset compared to the other two. This is because the negative feedback in KuaiRand is characterized by users’ explicit “dislike” actions. Compared to the “low-rated” and “exposure & unclick” types, “dislike” actions are more intense and targeted, which may result in slightly poorer performance of the editing methods.
5.3. Editing Efficiency (RQ2)
@@
An interesting phenomenon is that time costs for LightGCN+EFT and XsimGCN+EFT are lower than those for MF+EFT. This can be attributed to the insufficient model performance of MF. For LightGCN and XSimGCL, they import explicit collaborative information via graph, rendering the learned embedding distribution more compact. Therefore, the editing method only needs minor modifications to substantially change the target item’s order in the candidate set, whereas MF requires more significant modifications.
5.4. Explicit Editing Pairs Quantity (RQ3)
In real-world recommendation editing scenarios, the number of explicit editing pairs encountered is often uncertain, and their quantity can significantly impact model editing methods. We explore the impact of the number of explicit editing pairs on the performance metrics of model editing. Figure 3 illustrates how the ES, EC, EP, and EA metrics change for different methods as the number of explicit editing pairs varies, specifically for editing the XSimGCL model trained on the QB-video dataset. From the perspective of ES, the EFT method consistently maintains the best editing capability. Existing well-performing regularization-based and replay-based approaches also remain stable in terms of ES, while poorly performing methods like SML and EGNN see further declines in performance as the number of explicit editing pairs increases. From the perspectives of EC and EP, the majority of methods show a trend of increasing EC and decreasing EP with more explicit editing pairs. This is because more explicit editing pairs can provide more collaborative information but also increase the difficulty of editing, leading to a sacrifice in EP performance. For EA, EFT maintains 100% accuracy, whereas other methods decline as editing pairs increase. The L2 method uniquely shows opposite trends in EC and EP and suffers the most in EA, likely due to its overly stringent constraints.
5.5. Editing Objective (RQ4)
@@
Figure 4. The influence of proper objective.
6. Conclusion
@@ Appendix A Experimental Setting
A.1. Collaborative filtering model
@@
MF is a cornerstone in collaborative filtering for recommendation systems, decomposing a user-item interaction matrix into lower-dimensional latent factor matrices for users and items. This approach uncovers hidden patterns in data, enabling predictions of unobserved interactions by capturing underlying user preferences and item characteristics.
LightGCN simplifies the design of Graph Convolutional Networks (GCNs) for collaborative filtering in recommendation systems. It streamlines GCNs by removing feature transformation and nonlinear activation, focusing on the essential component of neighborhood aggregation to learn user and item embeddings. This process effectively captures the high-order connectivity in the user-item interaction graph, enhancing recommendation quality by leveraging the structural information within the graph.
XSimGCL is an improvement upon SimGCL, introducing a highly straightforward graph contrastive learning approach. Not only does it inherit the noise-based augmentation from SimGCL, but it also significantly reduces computational complexity by simplifying the propagation process.
A.2. Model Parameter Settings
@@
Figure 5. Editing efficiency.
Table 3. Experimental results (Mean ± Std.) of all editing methods.
Epinions MF LightGCN XSimGCL
Method EC EP Recall NDCG EC EP Recall NDCG EC EP Recall NDCG
FT 0.6485±0.0236 0.8224±0.0365 0.1312±0.0008 0.0607±0.0003 0.4580±0.0183 0.7879±0.0531 0.1564±0.0010 0.0741±0.0006 0.5932±0.0172 0.7276±0.0504 0.1692±0.0019 0.0835±0.0007
LWF 0.6554±0.0284 0.8080±0.0429 0.1308±0.0010 0.0605±0.0005 0.4494±0.0174 0.8094±0.0465 0.1569±0.0010 0.0743±0.0005 0.5949±0.0223 0.7250±0.0496 0.1688±0.0026 0.0834±0.0010
L2 0.6610±0.0247 0.8001±0.0354 0.1306±0.0006 0.0605±0.0003 0.4895±0.0117 0.8702±0.0183 0.1540±0.0006 0.0734±0.0004 0.5054±0.0213 0.9481±0.0141 0.1737±0.0004 0.0851±0.0001
SRIU 0.6564±0.0268 0.7997±0.0457 0.1306±0.0012 0.0604±0.0006 0.4498±0.0134 0.7612±0.0556 0.1570±0.0013 0.0741±0.0006 0.6068±0.0208 0.7073±0.0511 0.1680±0.0019 0.0829±0.0010
RSR 0.7936±0.0117 0.8823±0.0338 0.1301±0.0014 0.0595±0.0009 0.5800±0.0190 0.5236±0.0876 0.1367±0.0071 0.0653±0.0033 0.5853±0.0147 0.7376±0.0488 0.1691±0.0019 0.0835±0.0009
SPMF 0.7976±0.0123 0.8919±0.0243 0.1300±0.0011 0.0596±0.0007 0.5768±0.0179 0.5338±0.0916 0.1373±0.0070 0.0657±0.0033 0.5932±0.0159 0.7411±0.0495 0.1696±0.0019 0.0836±0.0010
SML 0.8230±0.0001 0.0010±0.0000 0.0021±0.0000 0.0008±0.0000 0.6907±0.0001 0.1216±0.0000 0.0776±0.0000 0.0355±0.0000 0.6983±0.0000 0.1036±0.0000 0.0779±0.0000 0.0356±0.0000
SiReN 0.5359±0.0331 0.6265±0.1316 0.1228±0.0068 0.0566±0.0036 0.6135±0.0771 0.2930±0.1859 0.0945±0.0377 0.0445±0.0182 0.5406±0.0491 0.4914±0.1653 0.1479±0.0203 0.0715±0.0100
SiGRec 0.0007±0.0003 0.9342±0.0297 0.1304±0.0015 0.0602±0.0008 0.6682±0.0612 0.2772±0.1175 0.1036±0.0230 0.0490±0.0109 0.6754±0.0345 0.4770±0.1080 0.1484±0.0120 0.0738±0.0057
EGNN 0.5629±0.0256 0.5293±0.0943 0.1181±0.0059 0.0548±0.0028 0.5866±0.0265 0.5082±0.0687 0.1389±0.0062 0.0662±0.0027 0.5455±0.0557 0.5361±0.1354 0.1556±0.0138 0.0761±0.0075
BiEGNN 0.6495±0.0275 0.8182±0.0387 0.1311±0.0008 0.0607±0.0004 0.4470±0.0239 0.8318±0.0194 0.1573±0.0008 0.0745±0.0003 0.5968±0.0199 0.7178±0.0593 0.1685±0.0028 0.0832±0.0013
EFT 0.6502±0.0390 0.8193±0.0482 0.1310±0.0008 0.0606±0.0003 0.6920±0.0200 0.8296±0.0444 0.1563±0.0011 0.0741±0.0005 0.5813±0.0368 0.8906±0.0271 0.1726±0.0006 0.0847±0.0002
KuaiRand MF LightGCN XSimGCL
Method EC EP Recall NDCG EC EP Recall NDCG EC EP Recall NDCG
FT 0.5355±0.0333 0.7172±0.0131 0.1764±0.0015 0.0656±0.0006 0.4421±0.0135 0.7882±0.0201 0.2329±0.0013 0.0875±0.0006 0.4265±0.0166 0.8300±0.0331 0.2494±0.0018 0.0956±0.0006
LWF 0.5350±0.0342 0.7189±0.0113 0.1766±0.0018 0.0657±0.0006 0.4428±0.0107 0.7806±0.0387 0.2328±0.0012 0.0877±0.0008 0.4287±0.0147 0.8296±0.0366 0.2493±0.0016 0.0956±0.0006
L2 0.5356±0.0335 0.7177±0.0129 0.1759±0.0019 0.0654±0.0008 0.4129±0.0255 0.9248±0.0235 0.2325±0.0016 0.0880±0.0004 0.4490±0.0167 0.7057±0.0192 0.2458±0.0023 0.0941±0.0008
SRIU 0.5350±0.0326 0.7174±0.0145 0.1764±0.0013 0.0656±0.0005 0.4401±0.0186 0.7467±0.0432 0.2326±0.0014 0.0875±0.0008 0.4314±0.0210 0.7908±0.0435 0.2492±0.0023 0.0952±0.0006
RSR 0.6493±0.0362 0.8702±0.0327 0.1808±0.0013 0.0670±0.0006 0.4831±0.0169 0.5132±0.0592 0.2033±0.0079 0.0768±0.0029 0.4367±0.0184 0.8237±0.0370 0.2486±0.0020 0.0954±0.0006
SPMF 0.6786±0.0377 0.8836±0.0323 0.1804±0.0012 0.0668±0.0004 0.4915±0.0161 0.4865±0.0656 0.1996±0.0084 0.0754±0.0033 0.4361±0.0187 0.8253±0.0382 0.2489±0.0018 0.0956±0.0007
SML 0.6486±0.0003 0.0564±0.0008 0.0591±0.0008 0.0189±0.0002 0.6340±0.0000 0.1871±0.0000 0.1649±0.0000 0.0613±0.0000 0.6279±0.0000 0.1756±0.0000 0.1646±0.0000 0.0611±0.0000
SiReN 0.4834±0.0316 0.5346±0.1299 0.1588±0.0129 0.0587±0.0050 0.5772±0.0818 0.3565±0.1685 0.1700±0.0492 0.0625±0.0198 0.6562±0.0441 0.2642±0.1340 0.1532±0.0482 0.0574±0.0188
SiGRec 0.0006±0.0002 0.9247±0.0271 0.1793±0.0012 0.0660±0.0007 0.6114±0.0476 0.2349±0.0772 0.1536±0.0219 0.0563±0.0091 0.5375±0.0456 0.4804±0.0953 0.2116±0.0157 0.0815±0.0063
EGNN 0.4682±0.0357 0.6083±0.1040 0.1695±0.0084 0.0631±0.0038 0.5236±0.0369 0.4398±0.0870 0.1947±0.0164 0.0736±0.0061 0.5119±0.0288 0.5119±0.0753 0.2234±0.0122 0.0857±0.0041
BiEGNN 0.5363±0.0334 0.7148±0.0165 0.1762±0.0017 0.0657±0.0006 0.4412±0.0097 0.7966±0.0316 0.2326±0.0015 0.0876±0.0009 0.4272±0.0165 0.8148±0.0361 0.2481±0.0019 0.0952±0.0005
EFT 0.5358±0.0338 0.7183±0.0142 0.1768±0.0016 0.0657±0.0005 0.5897±0.0212 0.8286±0.0275 0.2315±0.0013 0.0876±0.0004 0.5192±0.0237 0.8453±0.0109 0.2473±0.0004 0.0951±0.0002
QB-video MF LightGCN XSimGCL
Method EC EP Recall NDCG EC EP Recall NDCG EC EP Recall NDCG
FT 0.7370±0.0209 0.7779±0.0486 0.2893±0.0033 0.1748±0.0025 0.4829±0.0319 0.8023±0.0474 0.3182±0.0010 0.1965±0.0012 0.5028±0.0267 0.7803±0.0311 0.3462±0.0014 0.2148±0.0010
LWF 0.7390±0.0217 0.7657±0.0494 0.2888±0.0032 0.1747±0.0023 0.4773±0.0351 0.8263±0.0554 0.3185±0.0009 0.1968±0.0008 0.4910±0.0331 0.8085±0.0612 0.3477±0.0027 0.2159±0.0024
L2 0.7070±0.0131 0.8307±0.0282 0.2915±0.0019 0.1761±0.0015 0.4535±0.0187 0.9669±0.0070 0.3193±0.0002 0.1972±0.0002 0.4808±0.0270 0.8451±0.0696 0.3487±0.0019 0.2169±0.0014
SRIU 0.7406±0.0255 0.7620±0.0506 0.2887±0.0028 0.1744±0.0020 0.4651±0.0264 0.8139±0.0357 0.3185±0.0007 0.1970±0.0006 0.5448±0.0359 0.6865±0.0428 0.3383±0.0031 0.2100±0.0024
RSR 0.7367±0.0502 0.8794±0.0346 0.2910±0.0021 0.1731±0.0032 0.5750±0.0212 0.4657±0.0551 0.2688±0.0114 0.1605±0.0075 0.5207±0.0393 0.7081±0.0550 0.3397±0.0048 0.2109±0.0036
SPMF 0.8050±0.0152 0.8097±0.0181 0.2855±0.0017 0.1660±0.0024 0.5748±0.0295 0.4742±0.0846 0.2691±0.0168 0.1611±0.0112 0.5176±0.0312 0.7127±0.0341 0.3404±0.0031 0.2114±0.0021
SML 0.9000±0.0000 0.0001±0.0000 0.0003±0.0000 0.0002±0.0000 0.6962±0.0000 0.1376±0.0000 0.1686±0.0000 0.1015±0.0000 0.6870±0.0000 0.1105±0.0000 0.1648±0.0000 0.1000±0.0000
SiReN 0.5617±0.0449 0.4192±0.1517 0.2258±0.0464 0.1317±0.0311 0.6531±0.1108 0.3363±0.2369 0.1926±0.1093 0.1166±0.0678 0.6571±0.0849 0.3194±0.2502 0.2100±0.1086 0.1290±0.0690
SiGRec 0.0010±0.0004 0.9007±0.0464 0.2882±0.0034 0.1727±0.0027 0.6954±0.0604 0.3006±0.0811 0.2377±0.0266 0.1398±0.0173 0.5894±0.0372 0.5371±0.0376 0.3209±0.0099 0.1981±0.0080
EGNN 0.5243±0.0434 0.6291±0.0558 0.2767±0.0073 0.1652±0.0064 0.5722±0.0309 0.5001±0.0757 0.2814±0.0146 0.1713±0.0103 0.5628±0.0397 0.5231±0.0386 0.3062±0.0116 0.1890±0.0085
BiEGNN 0.7409±0.0231 0.7691±0.0300 0.2893±0.0017 0.1751±0.0011 0.4849±0.0224 0.7942±0.0289 0.3172±0.0021 0.1957±0.0021 0.4988±0.0383 0.7794±0.0843 0.3456±0.0038 0.2144±0.0028
EFT 0.7394±0.0245 0.7670±0.0429 0.2890±0.0024 0.1746±0.0018 0.6922±0.0156 0.8920±0.0255 0.3186±0.0005 0.1967±0.0006 0.5733±0.0195 0.9224±0.0043 0.3502±0.0003 0.2176±0.0002
A.3. Editing Method